OpenStudy (anonymous):

answer the function f (x)= x^2+3x When f=(x)-2

OpenStudy (shamil98):

when f(x) = x-2? or when f(x) = -2 ??

OpenStudy (anonymous):

well the way it is written it just says find f(x)-2

OpenStudy (anonymous):

\(f(x)-2 = x^2+3x-2\) but I don't think that's really what it's asking.

OpenStudy (anonymous):

? that is definitely not possible.

OpenStudy (anonymous):

Can you take a photo of the text?

OpenStudy (anonymous):

I'm guessing it's either \(f(x-2)\) or \(f(-2)\)

OpenStudy (shamil98):

yeah

OpenStudy (anonymous):

ill send a photo give me a second.

OpenStudy (anonymous):

i was told i was supposed to do something like f(x)-2=(x^2+3x)-2 but i dont understand that.

OpenStudy (anonymous):

Ok so... that's really all it was.

OpenStudy (anonymous):

f(x)-2 = x^2+3x-2

OpenStudy (anonymous):

is that it? i dont change any of it?

OpenStudy (anonymous):

Yes, that's really it.

OpenStudy (anonymous):

That's why nobody could believe that was what you were asking. it was too simple!

OpenStudy (anonymous):

okay. thank you!